Two equal forces (P each) act at a point inclined to each other at an angle of 120. The magnitude of their resultant is
ans.P

Grade:12th pass

1 Answers

Khimraj
5 years ago
Resultant of two vector
R  = \sqrt{A^{2} + B^{2} +2ABcos\Theta }
Then
 
R = \sqrt{P^{2} + P^{2} +2P^{2}cos120 }
R = \sqrt{P^{2} + P^{2} - P^{2} }
Then R = P
Hope it clears.
If you like answer then please approve it.